A chemist needs 9 liters of a 50% salt solution. All she has available is a 20% salt solution and a 70% salt solution. How much of each of the two solutions should she mix to obtain her desired solution? Show work please.
Answer by jorel1380(3719)

x+y=9
.20x+.70y=4.5
2x+7y=45
2x+2y=18
5y=27
y=27/5
x=18/5
She needs 18/5 liters 20% solution and 27/5 liters 70% solution..